THE PREDISSOCIATION MECHANISM FOR ^{2}Σ^{+} RYDBERG STATES OF CALCIUM MONOCHLORIDE

\underline{JASON O. CLEVENGER}, NICOLE A. HARRIS, and ROBERT W. FIELD, Department of Chemistry, Massachusetts Institute of Technology, Cambridge, MA 02139; JIAN LI, Department of Chemistry, Tsinghua University, Beijing, China 100084.

This talk summarizes experimental results from recent ion-dip spectroscopy studies of CaCl as well as previously unpublished optical-optical-double-resonance work with specific regard to predissociation processes of ^{2}Σ^{+} Rydberg states in the low-n* (n* < 7, IP-E ∼ 2500 cm^{−1}) region. A single repulsive state (assigned as ^{2}Σ^{+}) was found to be responsible for all observed predissociations of ^{2}Σ^{+} Rydberg states. The n*-dependent internuclear distances of the intersections between Rydberg and repulsive ^{2}Σ^{+} states were determined through the use of trial-and-error Franck-Condon calculations. Values of the n*-scaled electronic matrix elements governing the Rydberg ↔ repulsive state interaction were obtained from the measured linewidths (.6 < Γ < 1.2 cm^{−1}) and computed Franck-Condon densities. With the assumption of a oneparameter form for the repulsive curve, E(cm^{−1} ) = \frac{C_{12}}{R^{12}} + D_{e}, where C_{12} has the units Å^{12}cm^{−1} and D_{e} = 33171 cm^{−1} for CaCl, the optimum value of C_{12} was found to be 2.354 × 10^{8}Å^{12}cm^{−1}.